OpenStudy (adrimit):

what is the center of a 3-sphere passing through the points (0,0,0),(0,1,0),(1,0,0), and (0,1,1)?

OpenStudy (adrimit):

well a 2-sphere embedded in 3 dimensional Euclidean space

OpenStudy (phi):

yes, a 3-sphere needs 4 coords. You could put the center at (x,y,z) and define the vectors center - point i.e. (x,y,z) - (0,0,0) = (x,y,z) (x,y,z) - (1,0,0) = (x-1, y ,z) (x,y,z) - (0,1,0)= (x, y-1, z) (x,y,z) - (0,1,1) = (x, y-1, z)

OpenStudy (phi):

the length squared of each vector is the radius squared: A x^2 + y^2 + z^2 = r^2 B (x-1)^2 + y^2 + z^2 = r^2 C x^2 + (y-1)^2 + z^2 = r^2 D x^2 + (y-1)^2 + (z-1)^2 = r^2 equation A - eq B gives x^2 - (x-1)^2 =0 x^2 - (x^2 -2x +1) = 0 2x-1=0 x= ½

OpenStudy (phi):

you can find y and z using A-C and C-D or you can argue from symmetry. either way, we get y=½ , z= ½ radius will be \[ \frac{\sqrt{3}}{2} \] see attached figure.